OpenStudy (anonymous):

Find the equation (in slope-intercept form) of the line passing through the points with the given coordinates: (1,2), (3,-1)

OpenStudy (katrinakaif):

To find an equation, we must: 1) First find the slope 2) Use the equation y-y1 = m (x-x1) 3) Put it all together To find the slope of the given points, we use: y2-y1 ----- x2-x1 The given points are: (1,2) : (x1,y1) (3,-1) : (x2,y2) Now plug it in the equation above: y2 - y1 -1 - 2 -3 ------ = ------ = --- = 'm' x2-x1 3 - 1 2 SECOND STEP is to find to conclude to an equation using: (y-y1) = m (x-x1) Choose any one of those given points: In my case, lets choose (1,2) as (x1,y1) Now PLUG IN; y - y1 = m (x - x1) y - 2 = -3/2 (x - 2) y - 2 = -3/2x + 6/2 y - 2 = -3/2x + 3 y = -3/2x + 5 (Final Answer)

OpenStudy (katrinakaif):

I'd appreciate a 'Good Answer' :-) Danke

OpenStudy (anonymous):

Okay, I got the slope part, but how did you figure the second half?

OpenStudy (katrinakaif):

We first choose any one of the given points. In this case, I choose (1,2) to substitute in as (x1,y1) in the equation: (y-y1) = m (x-x1)